OpenStudy (anonymous):

find the exact (nondecimal) value of sin(13pi)

jimthompson5910 (jim_thompson5910):

Rule: sin(n*pi) = 0 where n is any integer

jimthompson5910 (jim_thompson5910):

the most common you are probably familiar with are sin(0*pi) = sin(0) = 0 sin(1*pi) = sin(pi) = 0 sin(2*pi) = sin(2pi) = 0